A vehicle of mass 1000kg is moving up an incline of 10° to the horizontal ground. determine the increase in potential energy when the vehicle covers 60m up the incline
1
Expert's answer
2020-09-14T10:22:32-0400

"PE=mgh=mgl\sin10°=1000\cdot 9.81\cdot60\cdot\sin10°=102209(J)"